If the weights of 600 students are normally distributed with mean of 50 kilograms and a variance of 16kilograms.



a.) Determine the percentage of students with weights lower than 55 kgs.



b.) How many students with weight between 50kgs and 55kgs.

1
Expert's answer
2022-05-16T23:36:45-0400

Let "X=" the weight of student: "X\sim N(\mu, \sigma^2)."

Given "\mu=50kg, \sigma=\sqrt{16}kg"

a)


"P(X<55)=P(Z<\dfrac{55-50}{\sqrt{16}})"

"=P(Z<1.25)=0.89435"

"0.89435(600)=537"

537 students


b)


"P(50<X<55)=P(Z<\dfrac{55-50}{\sqrt{16}})"

"-P(Z\le\dfrac{50-50}{\sqrt{16}})=P(Z<1.25)-P(Z\le 0)"

"=0.89435-0.5=0.39435"

"0.39435(600)=237"

237 students